OpenStudy (linda):

whats the cosine and sine of pi, abd the cosine and sine of 2pi over 3

OpenStudy (amistre64):

cos(pi) = -1 sin(pi) = 0 cos(2pi/3) = -1/2 sin(2pi/3) = sqrt(3)/2

OpenStudy (amistre64):

some parts of the unit circle